A solid conducting sphere of radius 30 cm has a charge of 4 uC. If the potential is zero atinfinity. Find the value of the potential at the following distances from the center of the sphere:a) 45 cmb) 30 cmc) 15 cm

1
Expert's answer
2021-04-16T07:24:59-0400

a) At r=45 cm from the center (r>R):


"V=\frac{kq}{r}=\frac{9·10^9·4·10^{-6}}{0.45}=79.9·10^3\text{ V}."

b) and c) At the surface and inside for "r\leq R" the sphere the field is zero and the potential is


"V=\frac{kq}{R}=\frac{9·10^9·4·10^{-6}}{0.3}=120·10^3\text{ V}."